If 15 men , woking 9 hours a day, can reap a field in 16 days , in how many days will 18 men reap the field , working 8 hours a day ?
Answer by ankor@dixie-net.com(22740)

If 15 men, working 9 hours a day, can reap a field in 16 days, in how many days will 18 men reap the field, working 8 hours a day?
:
Find the no. of man-hours required to do the field
15*9*16 = 2160 man hrs
:
let d = no. of day required with 18 men working 8 hrs a day
18*8*d = 2160
144d = 2160
d = 2160/144
d = 15 days
:
Check: 18 * 8 * 15 = 2160